OpenStudy (anonymous):

(x+2)^2-x^2=24 solve for x

myininaya (myininaya):

Multiply (x+2)(x+2)\[=x^2+2x+2x+4=x^2+4x+4\] So we have \[x^2+4x+4-x^2=24\] combine like terms! :)

myininaya (myininaya):

\[4x+4=24 \text{ since } x^2-x^2=0\]

OpenStudy (anonymous):

thanxxxxx

OpenStudy (anonymous):

myininaya sorry but where did you get the x for the 4?